For the IVP, dy/dx= f(x,y), y(x0)= y0, the continuity of f(x,y) and ∂f/∂y guarantees the unique solution of the problem.
True or false with full explanation
1
Expert's answer
2020-11-11T12:28:26-0500

Existence and Uniqueness of Solutions

Theorem: Let the function "f" and "\partial f/\partial y" be continuous in some rectangle "\alpha<x<\beta, \gamma<y<\delta," containing the point "(x_0, y_0)." Then, in some interval "x_0-n<x_0<x_0+h" contained in "\alpha<x<\beta" there is a unique solution "y=\varphi(x)" of the initial value problem


"dy\/dx=f(x, y), y(x_0)=y_0"

The statement is true.